OpenStudy (anonymous):

Anyone knw permutations?? How many permutations are possible with repeats for P(3,2)

OpenStudy (anonymous):

Can anyone explain how to do this?

OpenStudy (anonymous):

How you do that??

OpenStudy (anonymous):

I have this one too, can u explain how i do it? How many permuations are possible without repeats for P(5,4)

OpenStudy (anonymous):

what does "without repeats" mean? \[P(5,4)\] is a number

OpenStudy (anonymous):

it is \[5\times 4\times 3\times 2\]

OpenStudy (anonymous):

and \[P(3,2)\] is \(3\times 2\) not \(3^2\)

OpenStudy (anonymous):

So its 120? I dont understand like with repeats and without?

OpenStudy (anonymous):

What would p(5,4) be with repeats?

OpenStudy (anonymous):

me neither unless it is something your math teacher made up

OpenStudy (anonymous):

ok i think i get it

OpenStudy (anonymous):

with repeats i guess means you can repeat an element, it is not used up when you place it

OpenStudy (anonymous):

the notation for that is NOT \(P(5,4)\) but not matter, if you can repeat any of the 5 items four times, the answer is \(5^4\)

OpenStudy (anonymous):

also for the first one, i was wrong, it would be \(3^2\)

OpenStudy (anonymous):

Ok i got it, do you know combinations? How many combinations are possible without repeats for C(6,2) 10 20 12 15

OpenStudy (anonymous):

Do i just multiply 6x2?